OpenStudy (anonymous):

what is the value of y if (2,y) is a solution to the equation -3x+6y=-5

OpenStudy (compassionate):

-3(2) + 6y = -5 -6 + 6y = -15 6y = -15 - 6 6y = -21 y = -3.5

OpenStudy (precal):

 y=1/6